OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN!!! Which of the following is the conjugate of a complex number with 6 as the real part and -5 as the imaginary part? 6-5i -6-5i 6+5i -6+5i

OpenStudy (anonymous):

I've got no clue this time. Sry

OpenStudy (anonymous):

Its okayy. Can you tag some people who might?

OpenStudy (anonymous):

Umm

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

While I wait I have another question can you try to answer it?

OpenStudy (anonymous):

i dont know this im sorry

OpenStudy (anonymous):

Oh sry. Sure go ahead

OpenStudy (anonymous):

@LexiLuvv2431

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

Use the table below to answer this question: Table: x= -1, 0 ,2 Y= -3, -1, 3 Find the average rate of change for the given function from x = -1 to x = 2. -2 -1 1 2

OpenStudy (anonymous):

@Tide_Poole

OpenStudy (anonymous):

Oh sorry.

OpenStudy (anonymous):

I'm back now

OpenStudy (anonymous):

Do you know it?

OpenStudy (anonymous):

Oh that's slope. Rate of Change=slope

OpenStudy (anonymous):

It's asking for slope

OpenStudy (anonymous):

Yea I suck at slopee

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

It's positive one. It's moving to the right once every time

OpenStudy (anonymous):

Tada!

OpenStudy (anonymous):

Find the average rate of change for the given function from x = -1 to x = 2. -2 -1 Answer: 1 2

OpenStudy (anonymous):

Thank youuu

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

Glad to be of assistance!

OpenStudy (anonymous):

Heyy I have one more slope question if you dont mindd! @Tide_Poole

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

@LexiLuvv2431

OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (-2, -5) and (2, 3)? (y + 2) = -2(x + 5) (y - 2) = 2(x - 3) (y - 3) = 2(x - 2) (y + 3) = -2(x + 2)

OpenStudy (anonymous):

Do you know what the y-intercept is?

OpenStudy (anonymous):

@LexiLuvv2431

OpenStudy (anonymous):

Noo I don't

OpenStudy (anonymous):

Are you typing?

OpenStudy (anonymous):

Noo just waiting on your responsee. I'm trying to finish my exam and I need to get a good grade.

OpenStudy (anonymous):

Oh cus it looks like you're typing. At least it says you are anyway...

OpenStudy (anonymous):

Ohh idk why, I'm just waitingg

OpenStudy (anonymous):

Lol!

OpenStudy (anonymous):

Let me go get some graph paper real quick and I'll be right back. Okay?

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

Okay the y-intercept is -1 since that's where it crosses the Y-Axis

OpenStudy (anonymous):

Okay so then what would be the answer?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Thank youu

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

sry it took so long though

OpenStudy (anonymous):

Noo its okayy